OpenStudy (anonymous):

Danny made a mistake in the following problem. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer? Line 1 21 + 35 ÷ 7 + 6(2) Line 2 21 + 5 + 6(2) Line 3 21 + 11(2) Line 4 21 + 22 Line 5 43

OpenStudy (anonymous):

Ok, so for this we need to know PEMDAS for the order of operation problems. PEMDAS is the way we need to work these types of problems. First we do what's in parenthesis. Then we work with the exponents. After that, we work with multiplication or divison, whichever one comes first from left to right. Finally we do addition or subtraction, whichever one comes first from left to right. So we see that PEMDAS is: 1. Parenthesis, 2. Exponents, 3. Multiplication/Division, (whichever one comes first from left to right) 4. Addition/Subtraction (whichever one comes first from left to right)

OpenStudy (anonymous):

Line 3. 5+6(2) IS NOT EQUAL TO 11(2)

OpenStudy (anonymous):

Thank you lucky! :D

OpenStudy (anonymous):

Using this, we see that there is no parenthesis work (An example of this would be (5+7) or something like that. Then we go to the exponents, and we see there's none there. K so now we go to multiplication/divison: \[21+35\div7+6(2)\] We see that there is both division & multiplication. However we have to do whichever one comes first, so we do 35 divided by 7 which is 5. That then leaves us: \[21+5+6(2)\] we now do the 6(2) and that gives us: 21+5+12 and now we can finally add all this up which will give us 38 :)

OpenStudy (anonymous):

This means that line 3 is where the person made the mistake :) Hope it helped :)

OpenStudy (anonymous):

can you help me in another one ?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

Evaluate 3 − 24 ÷ 8 + 4 2

OpenStudy (anonymous):

hey question, what's between the 4 and the 2? :)

OpenStudy (anonymous):

oh nothing aha the 2 is supposed to be the square so 4 squared .-. sorry /.\ hehe

OpenStudy (anonymous):

oh ok np :) \[3-24\div8+4^2\]

OpenStudy (anonymous):

that's the problem then right? sry had to refresh page :)

OpenStudy (anonymous):

mine is freezing -_- and yes :)

OpenStudy (anonymous):

omg I hate when that happens, just want to throw my computer out the window and stomp all over it. Ok so we have to use PEMDAS again. -Parenthesis -Exponents -Multiplication/Division (going from left to right) -Addition/Subtraction (going from left to right) so we have: \[3-24\div8+4^2\] First, there's no parenthesis stuff so cross that off our list. Next, is exponents, and the \[4^2\] has an exponent, so we do that first and get: \[3-24\div8+16\] Now we do multiplication/division and we see that 24 divided by 8 fits into this so we do that and that gives us 3. Now we have: \[3-3+16\] we add/subtract this all up from Left to right, and get 3-3+16= 0+16= 16 and that's how you find that answer :)